OpenStudy (anonymous):

solve:3ln(5x)-ln(25x^2)=13

OpenStudy (anonymous1):

Start by using the following property: \[\ln a-\ln b=\ln\frac{a}{b}\]

OpenStudy (anonymous1):

Notice that \[3\ln(5x)=\ln((5x)^3)=\ln(125x^3)\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous1):

Could you solve it?

OpenStudy (anonymous):

still not sure

OpenStudy (anonymous1):

Could you wait a moment? I need to do something now and I'll be back in 10-20 min.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

hey i figured it out. thanks for pointing me in the right direction

OpenStudy (anonymous1):

OK, you're welcome.